Let n"\in"N consider the set of nxn symmetric matrices over R with the usual addition and multiplication by a scalar

3.1. Show that this set with the given operations is a vector subspace of Mnn

3.2. What is the dimension of this vector subspace?

3.3. Find a basis for the vector space of 2x2 symmetric matrices.


1
Expert's answer
2021-10-21T10:12:20-0400

1)

We depend on that "M_{n,n}" is a vector space with respect to usual matrix operation addition and multiplication by a scalar. To prove that space "S_{n,n}" of all symmetric matrices is a such vector space also is sufficient to prove that "S_{n,n}" is closed for mentioned operations. Let "A,B\in S_{n,n}" , so "\forall(i,j,1\le i,j \le n)(A_{i,j}=A_{j,i},B_{i,j}=B_{j,i})"

Let C=A+B,or "\forall(i,j,1\le i,j \le n)(C_{i,j}=A_{i,j}+B_{i,j})"

We must show that "C\in S_{n,n}" . Let i,j be any indexes such that "1\le i,j\le n"

Then "C_{j,i}=[by\space definition ]=A_{j,i}+B_{j,i}" But "A,B\in S_{n,n}" , therefore

"A_{j,i}=A_{i,j},B_{j,i}=B_{i,j}=>A_{j,i}+B_{j,i}=A_{i,j}+B_{i,j}=C_{i,j}"

Thus "C\in S_{n,n}"

Further, let A be any matrix from "S_{n,n}" and "c\in R" any number, our task is to prove that "(cA)\in S_{n,n}" or is symmetric. Let i,j- be any coefficients such, that "1\le i,j\le n" , then "(cA)_{i,j}=c\cdot A_{j,i}=[A\in S_{n,n}]=c\cdot A_{i,j}=(c\cdot A)_{i,j}" that is equivalent to "(cA)\in S_{n,n}"

2)

For linear space "S_{n,n}" basis may be formed from elementary matrices "E^{k,m}\in S_{n,n}" defined as "(E_{i,j}^{k,m}=1)\equiv((i=k,j=m)\lor (i=m,j=k))"

Basis in "S_{n,n}" is "\lbrace E^{k,m}, 1\le k\le m\le n \rbrace" with cardinalyty N=1+2+...+n="\frac{n\cdot (n+1)}{2}" and therefore "dim(S_{n,n})=\frac{n\cdot (n+1)}{2}" .

if "X=(x_{i,j})_{1\le i,j\le n}" any matrix from "S_{n,n}" then

"X=\sum_{1\le i \le j \le n}x_{i,j}\cdot E^{i,j}" - decomposition X by baisis.

3) If n=2 we have basis in "S_{2,2}=\lbrace \begin{pmatrix}
 1 & 0\\
 0 & 0
\end{pmatrix}.
\begin{pmatrix}
 0 & 1 \\
 1 & 0
\end{pmatrix},
\begin{pmatrix}
 0 & 0 \\
 0 & 1
\end{pmatrix}
\rbrace"